OpenStudy (anonymous):

does anybody know what 27 x 1/3^4 IS?

OpenStudy (chaise):

27 x 3^-4 = 1/3

OpenStudy (anonymous):

make both parts of the equation improper first then multiply

OpenStudy (anonymous):

27= 3^3 Add exponents 3^(3-4) =3^-1 or 1/3